an object initially has a velocity 6m/s, travel a distance of 24m with an acceleration of -4m/s. find the final velocity of the object .


1
Expert's answer
2021-11-16T13:40:49-0500

We can find the final velocity of the object from the kinematic equation (it seems there is a mistake in condition of the question, because the square root becomes imaginary, so let's suppose that the object travel a distance of 4 m):


"v^2=v_0^2+2ad,""v=\sqrt{v_0^2+2ad},""v=\sqrt{(6\ \dfrac{m}{s})^2+2\times(-4\ \dfrac{m}{s^2})\times4\ m}=2\ \dfrac{m}{s}."